Dublin University Rifle Club

Version Date : January 14, 2002 Filename : Status :

Risk Assessment

Mark Dennehy January 14, 2002

Executive Summary

Risk Assessment document for the Dublin University Rifle Club.

Executive Summary

Risk: Fire

Hazard  Consequence  Preventative Measure  Person
In Place  Needed  Responsible

Poor Electrical System Death/Serious Injury Smoke alarm

Buildings Office

replacement of fire extinguisher retardent (water with CO_{2})

Buildings Office

Smoking Fire No Smoking Policy Duty Range Offi-

cer

Buildings Office (Smoking, Gas storage, Corrosive chemical storage, Lumber storage)

Fire, Explosion, Gas leakage None None

Buildings Office

Risk: Personal Injury

Hazard  Consequence  Preventative Measure  Person
In Place  Needed  Responsible
Accident (tripping, etc)  Personal Injury  First-Aid Training  Chief Range Offi-

cer

Appropriate First-Aid Kit

DUCAC

Tripping Personal Injury None Buildings Office area to be kept clean

Buildings Office

Power washer to be stored

unobtrusively

Buildings Office

Small entrance door Tripping/Head injury None Change entrance door (allow wheelchair access at same time)

Buildings office, Department of Sport

Risk: Electrocution

Hazard  Consequence  Preventative Measure  Person
In Place  Needed  Responsible

Substandard wiring on and immediately outside building Electrocution None Rewiring

Buildings Office

Risk: Poisoning

Hazard  Consequence  Preventative Measure  Person
In Place  Needed  Responsible

Particulate Lead from range Respiratory disorders None

Improve ventilation Buildings Office

Tile powder/dust from Tilecutter Concrete/other dust from walls/rafters/buildings office Chemical fumes from Buildings Office Lead/Arsenic/Antimony ingestion from inadaquate hygiene after handling pellets/bullets Heavy Metal Poisoning and associated disorders Sink, Soap Hot Water tap Buildings Office

Risk: Legal Liability

Hazard  Consequence  Preventative Measure  Person
In Place  Needed  Responsible

Legal liability (for accident/injury claims or other legal actions)

Bankrupcy, Inability to recruit committee/range officers None Liability Insurance DURC Committee,

DUCAC, College

Risk: Chemical leakage

Hazard  Consequence  Preventative Measure  Person
In Place  Needed  Responsible

Chemical leaks from Buildings Office Damaged equipment, personal injury, respiratory disorders, poisoning None Unknown

Buildings Office

Risk: Sports Injury

Hazard  Consequence  Preventative Measure  Person
In Place  Needed  Responsible

Back/Neck/Wrist/Elbow strain, Repetitive Strain Injury, Chronic Nerve Damage Personal Injury (particularly for Junior shooters)

Higher-quality shooting

clothing, DURC Committee, DUCAC, College

Instructor/Coaching

courses, DURC Committee, DUCAC, College

More suitable equipment, DURC Committee, DUCAC, College

Student sports physiother-

apy service DURC Committee, DUCAC, College

Risk: Flooding

Hazard  Consequence  Preventative Measure  Person
In Place  Needed  Responsible

Flooding due to poor drainage Equipment, Range damage None Cleaning of drains in park lane

Buildings Office

Flooding due to leaking roof Equipment, Range damage None

Fix roof outside range Buildings Office

Risk: Hearing damage

Hazard  Consequence  Preventative Measure  Person
In Place  Needed  Responsible

Loud noise from rifles Permanent hearing impairment or loss Ear defenders (adaquately provided and mandatory for all members)

None Duty Range Officer

Executive Summary

Risk: Theft

Hazard  Consequence  Preventative Measure  Person
In Place  Needed  Responsible
Equipment theft  Equipment loss  Adaquate precautions

on away trips, secure storage, insurance DURC Committee, DUCAC/College (secure storage, insurance)

Security camera/lights for Park Lane

College

Risk: Equipment Malfunction

Hazard  Consequence  Preventative Measure  Person
In Place  Needed  Responsible

Equipment malfunction Personal Injury, Equipment damage/loss Regular maintainance, inspection, supervision None

DURC

Risk: Gunshot Wound

Hazard  Consequence  Preventative Measure  Person
In Place  Needed  Responsible

Ricochet

Personal Injury  Adaquate range design,  Buildings  Of-

fice/Department of Sport

Correct supervision Duty Range Offi-

cer

Accidental discharge Personal Injury Correct Supervision Duty Range Officer

Risk: Power Failure

Hazard  Consequence  Preventative Measure  Person
In Place  Needed  Responsible

No lighting Personal Injury Torches

DURC

Rewiring, Emergency lighting

Buildings Office

Accidental discharge of firearm Personal Injury, Ricochet Adaquate training of members Duty Range Officer